OpenStudy (anonymous):

in cyclic quadrilateral ABCD, the value of sinA/2 + sinB/2 - cosC/2 - cosD/2 =

OpenStudy (anonymous):

A+C=pi; B+D=pi. C=pi-A. cos(C/2)=cos(pi/2-A/2)=sin(A/2). cos(D/2)=sin(B/2). Thus the sum is zero.

OpenStudy (anonymous):

the value of 1/3*5+1/5*7+1/7*9....+1/23*25 is

OpenStudy (anonymous):

if(1-1/2)(1-1/3)(1-1/4)(1-1/5).......(1-1/19)(1-1/20)=x/20 ,then value of x

OpenStudy (anonymous):

1

OpenStudy (anonymous):

when the natural numbers 1,2,3 ---500 are written ,then the digit 3 appears x times .the value of x is

OpenStudy (anonymous):

if the difference between the simple interest and compound interest compounded annually on a certain sum for two years at 5% rate of interest is rupee 1,then the sum is

OpenStudy (anonymous):

in a quadrilateral ABCD ,if AB||CD,<D=2<B,AD=b and CD=a then the side AB is of length

OpenStudy (anonymous):

if x+1/x=m and x^2+1/x^3 =n ,then x^3+1/x^2is equals to

OpenStudy (anonymous):

the value of 1/3*5+1/5*7+1/7*9....+1/23*25 is